The mass of a helium atom is 6.66 × 10-27 kg. Compute the specific heat at constant volume for helium gas (in J/kg .K) from the molar heat capacity  at constant volume.

Grade:11

1 Answers

Navjyot Kalra
8 years ago
We have to find out the specific heat at constant volume for helium gas from the molar heat capacity at constant volume.
Since the molecular mass of helium is 4.00 g, therefore the specific heat at constant volume for helium gas would be,
(CV)H = (3/2 R) / (4.00 g/mol)
= (3/2 ×8.31 J/mol. K) / (4.00 g/mol)
= (3/2 ×8.31 J/mol. K) / ((4.00 g/mol) (10-3 kg/1 g))
= 3116 J/kg. K
Rounding off to three significant figure, the specific heat at constant volume for helium gas would be 3120 J/kg. K.